OpenStudy (anonymous):

According to the answer this function is invertible and they haven't specified any requirements on the function x(t). But surely x(t) can be constructed in a way that this function isn't invertible right?

OpenStudy (anonymous):

\[\int\limits_{-\infty}^t x(5\times\tau)d\tau\]